2(x-3)-3(x + 5)=3(x-2)-7

Answer by stanbon(75887)

2(x-3)-3(x + 5)=3(x-2)-7
2x-6-3x-15 = 3x-6-7
4x=-8
x=-2
Cheers,
Stan H.